Here's one question-


What is the total number of electrons in an atom with highest principal quantum number as 4?????????

Grade:11

3 Answers

Srinivas Rao
13 years ago

Such an atom shows electronic configuration ...4f145d16s2 and the differentiating electron enters ante penultimate 4f orbital. So, the atom is Lutecium with atomic number 71.

The last element of 4th period with heighest atomic number is with atomic number 36 is not the answer.

RAHUL RANJAN
13 years ago
32
VIBHOR GAUR
13 years ago

see frnd...........for principal quantum no. 4 the azimuthal quantum numbers r such:   

                                l=0(4s)

                                l=1(4p)

                                l=2(4d)

                                 l=3(4f)......so  heres a formula:NO. OF ELECTRONS IN A GIVEN SUB ENERGY LEVEL

                                                                      =2(2L+1)

FINDING IT::FOR l=0 SO  2(2*0+1)====2ELECTRONS(IN THE SAME WAY CALCULATE FOR OTHER SUBENERGY LEVELS   .....THEN ADD THEM THEY WILL BE LIKE 2+6+10+14=32)SO THERE R 32 TOTAL NO. OF ELECTRONS IN FULL PRINCIPAL QUANTUM NO.4)

hope u got it........!!!!!